The Dahlia Medical Center has

30

labor​ rooms,

19

combination labor and delivery​ rooms, and

55

delivery rooms. All of these facilities operate around the clock. Time spent in labor rooms varies from hours to​ days, with an average of about a day. The average uncomplicated delivery requires about one hour in the delivery room. The average time in a combination​ labor-delivery room is about 24 hours.

During an exceptionally busy​ three-day period,

103

healthy babies were born at Dahlia Medical Center.

66 babies were born in separate labor and delivery rooms and

37

were born in combined labor and delivery rooms. Which of the facilities​(labor rooms, combination labor and delivery​ rooms, or delivery​ rooms) had the greatest utilization​ rate?

The labor rooms OR

combination labor and delivery rooms OR

delivery rooms ?

had the highest utilization rate of

_____​%.

​(Enter your response as a percent rounded to two decimal​ places.)
Answer #1

We assume that the birth of babies is equally divided between separate Labour room and delivery rooms ( 33 each)

Labour rooms = 30 service time =24 hours, which means, mu= 1 customer /day

Babies born per day = 33/3 =11

Utilisation = 11/1x30 = 11/30 = 36.66%

Combined rooms =19 service time = 24 hours, mu = 1 customer /day

Babies born per day =37/3 =12.33

Utilisation = 12.33 /1x19 = 64.89%

Delivery rooms =55 service time =1 hour, mu = 24 customers /day

Babies born per day = 33/3 =11

11/55 x24 = 0.83%

Combined rooms have best utilisation rate of 64.89%